zarkam21:

Evaulate f(x)= 2x^2+3x+5 at f(x^2+x)

zarkam21:

SO the domain..... JUst need a refresher on how to determine the domain?

Vocaloid:

start by trying to find any x-values where the function is undefined the result from this calculation is 2 x^4 + 4 x^3 + 5 x^2 + 3 x + 5 are there any x-values that make this undefined?

zarkam21:

0?

zarkam21:

well we would just end up with zero

Vocaloid:

hm, not quite, if we plug in x = 0 we get 5 which is fine there aren't really any x-values that make this undefined so we say the domain is all real values

zarkam21:

not even neg

Vocaloid:

other ways to express this: (-infinity, infinity) or capital R with an extra bar (trying to find this)

Vocaloid:



Vocaloid:

there we go

Vocaloid:



zarkam21:

Vocaloid:

ah, yes since you have a square root the expression under the square root sign must be 0 or more so x^2 - 4 >= 0 x^2 >= 4 this is true for all values greater than/equal to 2 or less than/equal to -2

zarkam21:

I just need a clear answer lol not saying your wrong. I have my midterm in calc tomorrow so am just refreshing everything

Vocaloid:

the website is technically incorrect as it should be "all real values except [-2,2]" with closed brackets not parentheses

zarkam21:

so where are you getting the 2,-2 from

zarkam21:

is it because of x^2 is four and the square root is 2?

Vocaloid:

so the expression under the root sign is "x^2-4" we know that the expression under a root sign has to be 0 or more so we set up the inequality x^2 - 4 >=0 adding 4 to both sides x^2 >= 4 taking the square root of both sides tells us that x must be 2 or more, OR -2 or less

Vocaloid:

so instead of saying "what must x be" we can specify "what CAN'T x be" and since we can't have x-values between -2 and 2 we say "all real values EXCEPT [-2,2]"

Vocaloid:

or, you could also say (-infinity,-2] to [2,infinity) as an equally valid answer

zarkam21:

for the multiplication law of exponents ...ot os (x^a)^b=x^ab

zarkam21:

Do you leave it as x^ab or is it really x^a*b?

Vocaloid:

x^ab and x^a*b are technically the same thing the second one makes it more clear that you multiply a*b first before taking the exponent

zarkam21:

oh okay so you do multiply it

zarkam21:

the exponents?

Vocaloid:

multiply the exponents, yes

zarkam21:

okAY

Vocaloid:

so (x^2)^3 = x^6

zarkam21:

and what if it is the addition property like x^6x^7= it would obviously be x^13 but if the bases are not the same like x^6y^7?

zarkam21:

it would be like an error problem right?

Vocaloid:

x^6y^7 cannot be simplified further and must be left as is

zarkam21:

could you show me an example of distributive law over multiplication

Vocaloid:

FOIL is an example but for something simple A(B+C) = A*B + A*C

zarkam21:

Got it